The following financial information was summarized from the accounting records of Train Corporation for the current year ended December 31:

  Rails
   Division
  Locomotive
   Division
    Corporate
   Total
Cost of goods sold $45,100     $31,000    
Direct operating expenses 26,300     22,100    
Sales 95,400     66,000    
Interest expense $2,600    
General overhead 19,100    
Income tax 4,100    

The net income for Train Corporation is

a.$72,200

b.$11,100

c.$36,900

d.$85,300
Answer #1

Answer

  • Correct Answer = Option ‘B’ $ 11,100

Sales

$161,400

Cost of Goods Sold

$76,100

Gross Profits

$85,300

Direct operating expenses

$48,400

General Overhead

$19,100

Interest expense

$2,600

Income tax

$4,100

Total expenses

$74,200

Net Income

$11,100

--Working

Sales

=95400+66000

Cost of Goods Sold

=45100+31000

Gross Profits

=161400-76100

Direct operating expenses

=26300+22100

General Overhead

19100

Interest expense

2600

Income tax

4100

Total expenses

=48400+19100+2600+4100

Net Income

=85300-74200